mason06:

The following examples illustrate the inverse property of addition. Study the examples, then choose the statement that best describes the property.

yeyedontplay:

Please give us the possible options when writing a question? :D

mason06:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @yeyedontplay Please give us the possible options when writing a question? :D \(\color{#0cbb34}{\text{End of Quote}}\) there is no option

mason06:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @yeyedontplay Please give us the possible options when writing a question? :D \(\color{#0cbb34}{\text{End of Quote}}\) 5 + (-5) = 0 -1.33 + 1.33 = 0

yeyedontplay:

What is the statement that best describes the property?

mason06:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @yeyedontplay What is the statement that best describes the property? \(\color{#0cbb34}{\text{End of Quote}}\) The following examples illustrate the inverse property of addition. Study the examples, then choose the statement that best describes the property. 5 + (-5) = 0 -1.33 + 1.33 = 0 Inverse property of addition: For real numbers, a + = 0.